OpenStudy (anonymous):

Find the gradient of the tangent to the curve y=(3x-1)^3 (2x-1)^2 at the point where x=2

OpenStudy (anonymous):

take the derivative, probably using the product and chain rule then replace \(x\) by \(2\)

OpenStudy (anonymous):

you got the derivative, or you need help with it? no need so simplify using any algebra, because your job will then be to evaluate at \(x=2\) so you can do that no matter what form the derivative is in

OpenStudy (anonymous):

\[y=(3x-1)^3 (2x-1)^2 \] \[y'=3\times 3(3x-1)^2(2x-1)+2\times 2(x-1)(3x-1)^3\] by the product and chain rule just leave it, don't simplify replace \(x=2\) and do the arithmetic

OpenStudy (anonymous):

3(3 x-1)^2 2(2x-1)=450

OpenStudy (anonymous):

well maybe simplify to \[y'=9(3x-1)^2(2x-1)+4(x-1)(3x-1)^3\]

OpenStudy (anonymous):

hold on that is not the derivative you have a product use \((fg)'=f'g+g'f\) with \[f(x)=(3x-1)^3, f'(x)=9(x-1)^2, g(x)=(2x-1)^2, g'(x)=4(x-1)\]

OpenStudy (anonymous):

I am still not getting the answer

OpenStudy (anonymous):

Its 3525 but i am getting 1175

OpenStudy (anonymous):

\[(2 x-1)^2 (3 x-1)^3=108 x^5-216 x^4+171 x^3-67 x^2+13 x-1 \]The derivative of the RHS above is\[540 x^4-864 x^3+513 x^2-134 x+13 \]and when evaluated at x=2 is equal to 3525.